Two angles are supplementary. One angle is 28 degrees less than 3 times the other angle. Find the angles.
Answer by longjonsilver(2297)

supplementary angles add to 180 dgrees.

let first angle = x
second angle is 3x-28

so, x + 3x-28 = 180
4x-28 = 180
4x = 208
x = 208/4
x = 52

and second angle is therefore 128.

jon